Errata for

THE CUT LOCUS AND THE DIASTASIS OF A HERMITIAN SYMMETRIC SPACE OF COMPACT TYPE

(Osaka J. Math. 22 (1985), 863–870)

Hiroyuki Tasaki

October 13, 1999

p.865 Line 2 of Lemma 2: “ahd D_{i}” should read “and D_{i}”